A hitter during batting practice hits a ball at 167.2 km/h at an angle of 57.2 degrees to
the horizontal. How high does the ball go?
1
Expert's answer
2020-12-04T07:33:52-0500

Let's write the equation of motion of the ball in vertical direction:


"y=v_0tsin\theta-\dfrac{1}{2}gt^2, (1)"

here, "v_0 = 167.2\ \dfrac{km}{h}" is the initial velocity of the ball, "t" is the time of flight of the ball, "\theta=57.2^{\circ}" is the launch angle, "y" is the vertical displacement of the ball (or the height)  and "g=9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

Let's first find the time that the ball takes to reach the maximum height from the kinematic equation:


"v_y=v_0sin\theta-gt_{rise},""0=v_0sin\theta-gt_{rise},""t_{rise}=\dfrac{v_0sin\theta}{g}."

Then, we can substitute "t_{rise}" into the first equation and find the maximum height:


"y_{max}=v_0sin\theta\cdot\dfrac{v_0sin\theta}{g}-\dfrac{1}{2}g(\dfrac{v_0sin\theta}{g})^2,""y_{max}=\dfrac{v_0^2sin^2\theta}{2g},""y_{max}=\dfrac{(167.2\ \dfrac{km}{h}\cdot \dfrac{1000\ m}{1\ km}\cdot \dfrac{1\ h}{3600\ s})^2\cdot sin^257.2^{\circ}}{2\cdot 9.8\ \dfrac{m}{s^2}}=77.76\ m."

Answer:

"y_{max}=77.76\ m."